​Recently, More Money 4U offered an annuity that pays 5.4% compounded monthly. If $1,251 is deposited into this annuity every​ month, how much is in the account after 8 ​years? How much of this is​ interest?
Type the amount in the​ account: ​$nothing
​(Round to the nearest​ dollar.)
Type the amount of interest​ earned: ​$nothing
​(Round to the nearest​ dollar.)
1
Expert's answer
2020-09-28T20:59:55-0400

future value calculated by the formula

"FV=PMT\frac{(1+i)^n-1}{i}"

where FV future value

PMT = $1,251 payment every month

i = 5.4%/12 = 0.45% = 0.0045

n = 12*8 = 96 payment period in months


"FV=1251\frac{(1+0.0045)^{96}-1}{0.0045}\approx149798"


"FV" = $149,798 amount in the account


amount of interest​ earned

"FV -nPMT= 149798-96*1251=29702"

amount of interest​ earned $29,702


Answer: $149,798 amount in the account; amount of interest​ earned $29,702